yallbabynene08:

(10 - 7) - (14 - 12) is it 1

carmelle:

Correct it would be 1. 10 - 7 = 3 14 - 12 = 2 3 - 2 = ?

carmelle:

You can use the calculator option to help u with these types of questions. :)

yallbabynene08:

ok

carmelle:

yea